OpenStudy (anonymous):

Using l'hopital's rule: lim x→∞ [x^2 sin(1/x)]?

OpenStudy (anonymous):

Do you know what l'hopital's rule is?

OpenStudy (anonymous):

yes , i arrange the function to sin 1/x over 1/x^-2 to get 0/0

OpenStudy (anonymous):

but then after that, i take the derivative of both of them , but end up with 0 / infinity

OpenStudy (kainui):

You can do L'hopital's rule multiple times until you get an answer, assuming what you're doing is right to begin with.

OpenStudy (anonymous):

but i must get 0/0 or infinity/infinity to continue

OpenStudy (anonymous):

Just take the derivative of the original function, and see what you get when x goes to infinity.

OpenStudy (kainui):

What's 0/infinity equal to? Is it 0?

OpenStudy (anonymous):

i did , and the limit for that derivative is 0 over infinity

OpenStudy (anonymous):

That would be 0, but that wouldn't be the correct answer.

OpenStudy (anonymous):

yea..i saw the answer is 1

OpenStudy (anonymous):

for the derivative of sin (1/x), isnt it -1/x^2 (cos(1/x))?

OpenStudy (anonymous):

then this limit will be 0

OpenStudy (kainui):

Well I looked at it like this: \[\frac{ \sin x^{-1}}{ x^{-2} }\]\[\frac{ -x^{-2}\cos x^{-1}}{ -2x^{-3} }=\frac{ x }{ 2 }cos\frac{ 1 }{ x }\] And we can see that if we take x to infinity here we have infinity divided by 2 which is infinity and cos of 0 which is 1 times infinity. Infinity is the answer.

OpenStudy (anonymous):

oh , but isn't x^2 does not equal to x^-2?

OpenStudy (anonymous):

it should be 1/x^-2 right

OpenStudy (kainui):

If it's x^2 then it's equal to 1/(x^-2) is it not?

OpenStudy (anonymous):

ohhhh..i think i caught my mistake

OpenStudy (anonymous):

thanks! i put 1/ 1/ x^-2

OpenStudy (kainui):

Yeah it's like making a double negative lol, no problem, glad I could help.